Outback Outfitters sells recreational equipment. One of the company’s products, a small camp stove, sells for $110 per unit. Variable expenses are $77 per stove, and fixed expenses associated with the stove total $145,200 per month.

Required:

1. What is the break-even point in unit sales and in dollar sales?

2. If the variable expenses per stove increase as a percentage of the selling price, will it result in a higher or a lower break-even point? (Assume that the fixed expenses remain unchanged.)

3. At present, the company is selling 15,000 stoves per month. The sales manager is convinced that a 10% reduction in the selling price would result in a 25% increase in monthly sales of stoves. Prepare two contribution format income statements, one under present operating conditions, and one as operations would appear after the proposed changes.

4. Refer to the data in Required 3. How many stoves would have to be sold at the new selling price to attain a target profit of $77,000 per month?
Answer #1

1) Break even unit = 145200/(110-77) = 4400 Units

Break even sales = 4400*110 = $484000

2) New selling price = 110+11 = 121; variable cost = 77*1.1 = 84.70

Break even unit = 145200/(121-84.70) = 4000 Units

So if variable expense increase as a percentage of selling price increase break even point will be lower.

3) Two contribution margin income statement

Present Proposed
Total Per unit Total Per unit
Sales 1650000 110 1856250 99
Variable cost 1155000 77 1443750 77
Contribution margin 495000 33 412500 22
Fixed cost 145200 145200
Net income 349800 267300

4) Required sales = (145200+77000)/22 = 10100 Units